OpenStudy (anonymous):

Eight less than a number doubled is twenty more than the number. What is the number? A. 28 B. 6 C. 14 D. 36

OpenStudy (anonymous):

I just need the equation other wise I can figure it out on my own

OpenStudy (anonymous):

Ok so what you need to do is to substitute the numbers

OpenStudy (anonymous):

8-x^2= 20+x

OpenStudy (anonymous):

?

jigglypuff314 (jigglypuff314):

I think the "doubled" part should be times 2 and not ^2 because ^2 would have been "squared"

jigglypuff314 (jigglypuff314):

and "eight less than a" would be 2x - 8 and not 8 - 2x

OpenStudy (anonymous):

so 2x?

OpenStudy (anonymous):

ok ... so 2x-8=20+x?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

have you found the answer?

OpenStudy (anonymous):

yes it 28

OpenStudy (anonymous):

its*

OpenStudy (anonymous):

good work:)

OpenStudy (anonymous):

thanks :)